OpenStudy (anonymous):

Elaine is painting a fence. She knows that she needs 1 gallon of paint for each 10 feet of fencing. The fencing will enclose a rectangular yard that is 3 yards long and 2 yards wide. How much paint does Elaine need?

OpenStudy (anonymous):

P=2(L+W) P=2(3+2) P=10 yards Convert yards to feet multiply by 3 so 10 yards=30 feet 1/10=x/30 x=30*1/10=3 Gallons of paint Therefore Elaine needs 3 gallons of paint to paint the fence.